### v1

# OPRA-Vis: Visual Analytics System to Assist Organization-Public Relationship Assessment with Large Language Models

Abstract
Analysis of public opinions collected from digital media helps organizations maintain positive relationships with the public. Such public relations (PR) analysis often involves assessing opinions, for example, measuring how strongly people trust an organization. Pre-trained Large Language Models (LLMs) hold great promise for supporting Organization-Public Relationship Assessment (OPRA) because they can map unstructured public text to OPRA dimensions and articulate rationales through prompting. However, adapting LLMs for PR analysis typically requires fine-tuning on large labeled datasets, which is both labor-intensive and knowledge-intensive, making it difficult for PR researchers to apply these models. In this paper, we present OPRA-Vis, a visual analytics system that leverages LLMs for OPRA without requiring extensive labeled data. Our framework employs Chain-of-Thought prompting to guide LLMs in analyzing public opinion data by incorporating PR expertise directly into the reasoning process. Furthermore, OPRA-Vis provides visualizations that reveal the clues and reasoning paths used by LLMs, enabling users to explore, critique, and refine model decisions. We demonstrate the effectiveness of OPRA-Vis through two real-world use cases and evaluate it quantitatively, through comparisons with alternative LLMs and prompting strategies, and qualitatively, through assessments of usability, effectiveness, and expert feedback.
Index Terms:
Public relations, public opinion assessment, visualization, large language model, chain-of-thought prompting
I Introduction
Citizens actively express their thoughts and emotions about issues they face [1]. Social organizations, such as corporations and governments, influence individuals, while these individuals, including publics and stakeholders, also shape organizational actions [2, 3]. Their communicative behaviors generate opinion data [4, 5], reflecting experiences and revealing shifts in relationship quality. Understanding such data provides insights into public expectations, aiding organizations in strategy, policy, and legitimacy management [6].
Public relations (PR), a discipline of strategic management, relies on analyzing public expectations to maintain legitimacy and secure resources. Traditionally, researchers have used Organization-Public Relationship Assessment (OPRA) [7, 8] with Likert scale surveys to quantify the public’s relationship with organizations [4, 9]. However, these self-reports are prone to recall bias and social desirability, and cannot fully capture the dynamic nature of relationships due to their one-time format. To overcome these limitations, scholars are increasingly turning to social media and other online platforms, producing abundant textual data reflecting public opinion [4, 5]. Yet, interpreting such online texts poses linguistic challenges, such as jargon, quotations, and missing context, and requires converting unstructured input into OPRA’s structured dimensions. These difficulties require adopting Natural Language Processing (NLP) techniques to analyze text on a scale while preserving theoretical alignment with PR constructs.
Large Language Models (LLMs) support diverse NLP tasks and are widely adopted through pre-trained language models (PLMs) [10, 11]. While PLMs reduce the burden of training from scratch, integrating PR theory into LLMs for public opinion analysis faces three challenges. C1: LLMs require large labeled datasets for optimal performance on new tasks [10, 12]. C2: Labeling based on PR theory demands labor-intensive human coding [1, 13]. C3: LLMs provide limited transparency in how they map unstructured text to OPRA dimensions, making it challenging to confirm or adjust the rationale behind their relational judgments [14, 15, 16].
To tackle the three challenges (C1–C3), we introduce OPRA-Vis, a novel visual analytics (VA) system that integrates OPRA concepts of PR, LLMs, and visual analysis. The assessment and reasoning of PR experts is replicated by the system through the use of Gemma [17] and Chain-of-Thought (CoT) prompting with clues and reasons [15], based on few-shot examples (FSEs) [18]. Bidirectional Encoder Representations from Transformers (BERT) [19] is used to classify the sentiment of words in sentences, thereby providing PR analysis with insights into the linguistic phenomena that LLMs comprehend. Visual analysis modules interpret and interact with LLM decision processes through scatter plots with tag clouds, attention visualization, and prompt-editing interactions. We evaluate OPRA-Vis using two real-world datasets, conduct quantitative comparisons with baseline classifiers, and provide qualitative assessments through expert feedback and usability tests. Our main contributions are summarized as follows.
  • We propose OPRA-Vis, an interactive visualization integrating PR theory (e.g., OPRA) and LLMs for labeling.
  • We present an OPRA-informed continuous encoding and use a gravity model to visualize multiple concepts.
  • We allow users to adjust LLM-generated labels by exposing the clues and reasoning behind the labeling procedure.
All supplementary materials, including the demonstration video, source codes, and LLM prompts used in OPRA-Vis, are available at https://github.com/DVL-Sejong/OPRA-Vis.
II Related Work
In this section, we review the underlying techniques, including LLMs, VA for reasoning, and PR.
II-A Large Language Models (LLMs)
Various state-of-the-art (SOTA) pre-trained LLMs, such as BERT [19] and Generative Pre-trained Transformers (GPT) [20], have demonstrated effectiveness across a wide range of NLP tasks. These models are typically fine-tuned for task-specific applications while retaining their original architectures. For example, SCIBERT [10] adapts BERT to scientific domains. To reduce fine-tuning costs, ULMFiT [12] introduced transfer learning for NLP, though its generality is limited. Prompting methods, such as in-context learning [21, 22], guide LLMs via task descriptions and examples, enabling few- or zero-shot learning. CoT prompting [23, 24] improves by incorporating reasoning steps into the prompt, but remains highly sensitive to prompt design [25].
LLMs increasingly automate text classification [11, 26], aiming to reduce human labeling efforts. Meng et al. [27] proposed self-training without labeled data, while Wang et al. [14] combined GPT-generated pseudo labels with human annotations. However, in domain-specific PR analysis, in-context learning still faces reliability challenges. Our approach differs by integrating expert-guided CoT prompting with interactive visualizations that expose model rationales and support prompt and label refinement for OPRA.
II-B Visual Analytics for Reasoning from Text Data
VA facilitates effective decision-making by offering insights into data interpretation and human knowledge transformation [28]. Tag clouds are a widely used method for identifying topics based on word frequency [29]. Knittel et al. [30] introduced an interactive tag cloud visualization that captures both temporal and semantic dynamics. Such approaches have been applied to text labeling analysis [31, 32]. For example, Khayat et al. [31] developed VASSL, a system for labeling social spambots—automated entities that influence human behavior—by combining dimensionality reduction, sentiment analysis, and topic modeling. VASSL incorporates tag clouds for topic analysis and provides controls over dimensionality reduction parameters to detect malicious behavior. Similarly, Li et al. [32] proposed EvoVis for multi-class labeling, enabling users to explore relationships over historical overviews and supporting iterative analysis of labeling tasks.
Interactive VA systems also facilitate the interpretation of model decisions [33, 34, 35]. BertViz [33] offers multiscale visualizations of transformer attention heads, neurons, and model structures, although it does not support interactive editing. AttentionViz [34] visualizes joint query-key embeddings to reveal attention trends, but requires substantial pre-computation for large models. KnowledgeVIS [35] compares fill-in-the-blank prompts by generating semantic variations and interactive views, yet it has limitations in capturing fine-grained linguistic nuances. PrompTHis [36] investigates how editing prompts influences image generation by highlighting word-level changes that affect the outputs.
These studies use LLMs and prompts to examine how input prompts affect model behavior. In contrast, our approach leverages prompting to overcome the scarcity of labeled data. Specifically, we utilize expert instructions combined with CoT prompting that includes clues and reasons to embed domain knowledge into LLMs. Additionally, we visualize how these clues influence the model’s reasoning for labeling.
II-C Public Relations (PR)
The industry increasingly adopts digital media monitoring to track public sentiments and opinions toward organizations. Tools like Brandwatch [37] and other social media monitoring platforms support this process [38, 39]. Academic researchers also use social network analysis to examine the evolution of public opinion. For instance, Himelboim et al. [40] analyzed Twitter followership to identify social mediators linking the U.S. Department of State with international publics, revealing that informal actors primarily mediated relationships in the Middle East and North Africa.
PR researchers also study public opinion on politics and government [1, 41, 42, 43]. Political communication analysis in PR involves supporting initiatives of political parties, governments, and civil organizations through communication strategies [41]. It also helps mitigate public concerns on social issues [43]. Chon and Park [43] examined OPR during disease outbreaks, aligning behavioral intentions with the situational theory of problem solving [1] to promote Centers for Disease Control and Prevention (CDC) guideline adherence. PR studies further suggest that governments have a greater experience in crisis handling than citizens [42]. Lee and Jun [42] provided a theoretical view of public diplomacy, explaining how the U.S. and Korea foster cooperative relationships through international public experiences. This analysis also extends to public opinions from areas such as volunteerism [44].
Despite growing interest in artificial intelligence (AI), academic research applying deep learning or LLMs in PR remains scarce. Existing studies rely primarily on content analysis [45, 46] or case studies [47]. A key obstacle is the lack of labeled data required to train such large AI models. In PR, expert-driven human coding remains essential. To address this challenge, we adopt LLMs and CoT prompting with clues and reasons to emulate expert reasoning by using structured FSEs for assessing public opinion data.
III Design Considerations
This section presents the definitions of OPRA concepts, expert requirements, and the design goals of OPRA-Vis.
III-A Organization-Public Relationship Assessment (OPRA)
The OPRA [7] is a framework devised by scholars in PR to evaluate relationships, featuring four fundamental concepts [13]: Trust, Satisfaction, Commitment, and Control Mutuality. Trust comprises competence, dependability, and integrity. Satisfaction relates to positive emotional responses. Commitment concerns long-term expectations, while Control Mutuality addresses perceptions of balanced control. OPRA provides metrics for gauging these relationships, offering valuable insight to maintain favorable public connections. These metrics are typically obtained via surveys, supporting relationship improvement strategies across sectors such as business and government, and helping organizations strengthen public engagement over both short- and long-term horizons. PR analysis involves two distinct classification tasks: the binary assessment of OPRA concepts (labeling opinions as True or False for each concept) and sentiment analysis, which classifies the emotional tone of public opinion as Positive or Negative.
III-B Expert Requirements
Labeling public opinion data with OPRA concepts requires human coding by PR experts, which is resource-intensive. To mitigate this burden, we conducted semi-structured meetings with two PR experts. We synthesized a non-exhaustive but sufficient set of requirements to guide our design, with both experts later joining the project as coauthors.
  • R1
    PR experts lack labeled data and therefore need techniques that do not rely on training datasets.
  • R2
    They need a labeling tool that uses PR expertise to replace the inconsistent and labor-intensive human coding.
  • R3
    They aim to utilize OPRA concepts and sentiment for a detailed examination of ambiguous cases.
  • R4
    They require information that allows them to assess whether the model’s reasoning for labeling is reasonable.
  • R5
    They require simple, intuitive interactions to correct labeling errors and refine the reasoning for labeling.
  • R6
    They need a system that incorporates their edits across the remaining OPRA tasks to save time and effort.
III-C Design Goals
The system design goals for the conjunction use of PR and LLM are as follows.
  • G1
    LLMs leverage PR theory using a small number of exemplars (C1 & C2 in Section I, R1, R2 in Section III-B).
  • G2
    To contextualize ambiguous cases, the system quantifies the certainty of OPRA-related clues in the text and classifies the sentiment of sentences (C2, C3, R3).
  • G3
    The system presents the LLM’s reasoning and decision process (C3, R4).
  • G4
    Interactions are needed for PR experts to correct LLM decisions and errors (C2, C3, R5).
  • G5
    The system enhances usability through task-relevant filtering informed by OPRA labeling (C2, C3, R6).
Integrating PR theory with LLMs for labeling public opinion data faces three main challenges (C1–C3) as introduced in Section I. To address the challenge of insufficient training data for fine-tuning LLMs (G1), we implement the Gemma [17] model with CoT prompting, based on in-context learning. This method leverages textitFSEs provided by PR experts to label the data. Moreover, we have developed an LLM pipeline that represents OPRA-related clues in public opinion data as a continuous encoding (G2). Adopting a BERT [19] model reduces the dependence on human effort to classify sentiment by analyzing public opinion sentences. Sentiment classification enriches labeled data and broadens the analyses supported by our system (G2).
We combine visual analytics with CoT prompting that incorporates clues and reasons to audit and revise the model’s reasoning for labeling, supported by coordinated views and prompt editing (G3, G4, G5). CoT prompting elicits “think-aloud” rationales, making the LLM’s reasoning for labeling explicit and auditable (G3). Additionally, understanding the transformer’s attention mechanism is crucial for comprehending the LLM’s reasoning for labeling and for guiding prompt editing in CoT-based revisions. Therefore, attention visualization is devised to provide an attention mechanism for the transformer that influences LLM’s labeling (G3, G4). Moreover, public opinion data, i.e., freeform text posted on digital media, is diverse and noisy. Therefore, techniques that support usability are required for labeling, reasoning, and editing. We adopt sentence embeddings and dimensionality reduction to encode the certainty relevant to the OPRA concept (G2). This encoded certainty is then visualized in a scatter plot, with filtering applied based on certainty levels. The continuous encoding with certainty-based filtering for OPRA labeling improves usability and reduces human workload in visual analysis (G5).
Figure 1: Dataflow of OPRA-Vis. OPRA-Vis takes FSEs, expert instructions for each concept, and public opinion sentences as input and produces OPRA concept-labeled data. (a) shows initial concept labeling (True or False), (b) shows sentiment classification of sentences followed by frequency calculation, and (c) computes a certainty of concepts (CoCCoC). T., S., C., and C.M. are Trust, Satisfaction, Commitment, and Control Mutuality, respectively.
IV Data Processing of OPRA-Vis
[…]
IV-A Concept Labeling with Clues and Reasons
[…]
Figure 2: The system overview of OPRA-Vis. OPRA-Vis consists of three tightly coordinated views. (a) is an OPRA Concept Space view to explore the certainty of concepts ($CoC$) relying on OPRA concepts using a scatter plot and provide the sentiment of words in sentences using tag clouds. (a-1) shows a distribution chart displaying data point density along the selected concept axis. (b) is a Data Management view. The Data Management view features a slider for data filtering, as presented in (b-1). The filter uses the CoCCoC range to update point visibility in the scatter plot in (a). The data table includes certainty, concept labels from LLM and experts, and data exclusion, as shown in (b-2). (c) is an LLM Reasoning & Editing view for analyzing LLM’s reasoning for labeling and editing prompts to adapt the user’s knowledge. This view provides an attention visualization for analyzing the effectiveness of sentences in prompting. (c-1) is a pop-up window that organizes the generated sentences, focus sentences, and attention scores. The pop-up window is activated by hovering over the dot representing the sentence of attention visualization. (c-2) shows buttons to open the prompt editor and run the assessment using the edited prompt.
IV-B Sentiment Classification
[…]
IV-C Certainty of Concepts (CoC)
While the concept labels yield outputs of True or False, mirroring PR expert labeling, public opinion sentences often contain ambiguities that complicate classification into clear-cut categories. To reflect these nuances in analysis, we define a concept-level certainty (certainty of concepts; $CoC$) that measures how strongly OPRA clues in a sentence support a given OPRA concept, as shown in Figure 1 (c). Our system uses the $CoC$ to provide certainty-based filtering in the analysis of LLM’s reasoning for OPRA concept labeling.
To compute the $CoC$ for OPRA-informed clues, we embed sentences using ${\text{BERT}}_{\text{BASE}}$, as shown in Figure 1 (c). We apply dot-product attention between the sentence embedding and expert instructions of each OPRA concept label (True and False). Dot-product attention measures semantic similarity between embeddings in the same vector space, enabling direct comparison without introducing additional parameters. Let $\mathbf{h}_{i}\in\mathbb{R}^{d_{k}}$ be the query embedding of the $i$-th sentence and $\mathbf{C}_{j}^{(\ell,k)}\in\mathbb{R}^{d_{k}}$ be the $k$-th key embedding for concept $j$ with label $\ell\in\{\text{true},\text{false}\}$, where $d_{k}$ is the embedding dimension. We compute the certainty of concepts $CoC$ as follows.
$w_{j}^{(i,\ell)}$ $=\text{mean}\left(\max_{k}\left[\text{softmax}\left(\mathbf{h}_{i}\cdot\left(\mathbf{C}_{j}^{(\ell,k)}\right)^{T}/\sqrt{d_{k}}\right)\cdot\left|\mathbf{C}_{j}^{(\ell,k)}\right|\right]\right)$
$CoC_{ij}$ $=\begin{cases}w_{j}^{(i,\text{true})}/\sum_{\ell}w_{j}^{(i,\ell)},&\text{if }\ \sum_{\ell}w_{j}^{(i,\ell)}>0\\ 0.5,&\text{otherwise}\end{cases}$ (1)
Here, the max operation selects the highest attention score within each expert instruction, while the mean operation averages these maximum scores across multiple expert instructions for the same concept-label pair. The computed $CoC$ values are standardized and rescaled to $[0,1]$ for consistent comparison and intuitive interpretation: values near $1$ indicate high certainty for True OPRA concepts, values near 0 indicate high certainty for False OPRA concepts, and $0.5$ denotes neutral certainty. These $CoC$ values serve as quantitative indicators of concept certainty, supporting analytical evaluation.
V OPRA-Vis: LLM-Assisted Visual Analytics System with CoT Prompting for OPRA
This section presents OPRA-Vis, a system designed to assist PR experts in OPRA labeling through three interconnected components (Figure 2): (a) an OPRA Concept Space view, (b) a Data Management view, and (c) an LLM Reasoning & Editing view. These components are tightly integrated, with interactions between (a) and (b) as well as between (b) and (c), enabling seamless exploration, analysis, and iterative editing.
V-A OPRA Concept Space View
Figure 3: Position adjustment using the gravity model weighted by CoCCoC. (a) Before: Initial t-SNE layout of sentence embeddings forms dense clusters without clear concept alignment. (b) After: A gravity model uses True and False concept vertices ($\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ in Equation 2) as gravitational attractors, where each sentence is pulled toward its corresponding concept edge with force proportional to its certainty (CoCCoC). High-certainty sentences experience a stronger gravitational pull and move closer to concept edges. In contrast, low-certainty sentences experience weaker forces and remain near the center, resulting in eight radial structures separating concepts and certainty levels.
The OPRA Concept Space view, shown in Figure 2 (a), visualizes public opinion texts according to the four OPRA concepts—Trust, Satisfaction, Commitment, and Control Mutuality—using a scatter plot and tag clouds. We adopt an octagonal layout that uses the angle to indicate the direction of the concept-specific True or False opposing label and the radius to encode $CoC$ in Equation 1. Sentence embeddings are represented as points inside the octagon, while tag clouds are positioned around its perimeter. The embeddings are generated with $\text{BERT}_{\text{BASE}}$ [19], producing 768-dimensional vectors that are projected into 2D using t-SNE [48]. To better reflect OPRA-specific relationships, we further adjust the positions using Equation 2, which incorporates the $CoC$. The gravitational adjustment ensures that as $CoC$ increases, points move toward the corresponding concept edge. To achieve this adjustment, we employ a physics-inspired gravity model [49], where the vertices representing OPRA concepts act as gravitational attractors. Sentence positions are iteratively updated based on their $CoC$ values through a constrained simulation.
$\mathbf{v}_{\text{target}}^{(i,j)}$ $=\mathbf{v}_{j}^{+}\cdot\textbf{1}_{CoC_{ij}>0.5}+\mathbf{v}_{j}^{-}\cdot\mathbf{1}_{CoC_{ij}<0.5}$
$\alpha_{ij}$ $=|CoC_{ij}-0.5|\cdot\alpha_{\text{base}}\cdot\mathbf{1}_{CoC_{ij}\neq 0.5}$
$F_{ij}$ $=G\frac{\alpha_{ij}}{(r_{ij}+\epsilon_{1})^{2}};\quad\textbf{F}_{ij}=F_{ij}\cdot\frac{\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}}{\left|\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}\right|+\epsilon_{2}}$
$\mathbf{u}_{i}^{(t+1)}$ $=\gamma\cdot\mathbf{u}_{i}^{(t)}+\delta\cdot\sum_{j}\mathbf{F}_{ij};\quad\mathbf{p}_{i}^{(t+1)}=\mathbf{p}_{i}^{(t)}+\mathbf{u}_{i}^{(t+1)}$ (2)
Here, $\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ represent True and False concept vertices for concept $j$, $\mathbf{1}_{\{\cdot\}}$ is the indicator function, $\alpha_{\text{base}}$ is the base gravitational strength, $G$ is the gravitational constant, $r_{ij}=\|\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}\|_{2}$ is the Euclidean distance, $\mathbf{u}_{i}^{(t)}$ is the velocity vector, $\gamma$ is the damping factor, $\delta$ is the step size, $\epsilon_{1}$ prevents force divergence when $r_{ij}=0$, and $\epsilon_{2}$ prevents division by zero when computing directions. In our implementation, we use $\alpha_{\text{base}}=2.0$, $G=1.0$, $\gamma=0.8$, $\delta=0.1$, $\epsilon_{1}=0.01$, and $\epsilon_{2}=10^{-10}$. These parameters were empirically selected and can be adjusted. We will explore systematic optimization methods in future work. The gravitational force follows an inverse square law, where $\alpha_{ij}$ acts as the gravitational mass proportional to concept certainty. When $CoC_{ij}=0.5$, no gravitational force is applied as the sentence shows neutral certainty. Positions are constrained within a unit circle through normalization to maintain visual coherence. This iterative gravitational simulation continues for up to 200 iterations or until position changes fall below $10^{-4}$, creating spatial clustering where sentences are gravitationally attracted toward vertices corresponding to their respective concepts. Figure 3 illustrates this transformation: (a) shows the initial t-SNE layout, and (b) shows the adjusted plot reflecting $CoC$ values with gravitational attractors.
Figure 4: Interaction of the OPRA Concept Space view. Selecting an edge of the octagonal layout activates the corresponding OPRA concept; the scatter plot then color-encodes the CoCCoC (Section IV-C) for that concept to provide an overview of how opinions are distributed. (a) to (d) show the views with Trust, Satisfaction, Commitment, and Control Mutuality selected, respectively.
The scatter plot enables interactive concept selection: clicking on an OPRA concept label on the concept edge updates the color encoding based on the $CoC$ values for that concept. We use a blue-yellow-red color scheme, where blue indicates low certainty (near 0), yellow represents neutral certainty (around 0.5), and red corresponds to high certainty (near 1). This three-color gradient aligns naturally with the $CoC$ scale, which is centered around the neutral value of 0.5. Figures 4 (a)–(d) illustrate scatter plots for the cases when the user selects Trust, Satisfaction, Commitment, and Control Mutuality, respectively.
To highlight regions of high and low certainty, we provide a distribution chart for the user-selected OPRA concept label from the scatter plot, as illustrated in Figure 2 (a-1). The data are aligned along an axis whose endpoints correspond to the concept’s True and False values. Each point is orthogonally projected onto this axis, producing a position in $[0,1]$ based on its $CoC$. The distribution is shown as a histogram with a $\log_{10}$-scaled y-axis to emphasize patterns across frequency ranges. Users can switch between linear, $\ln$, $\log_{2}$, and $\log_{10}$ scales by clicking the chart title, allowing flexible analysis of both sparse and dense regions.
To facilitate analysis of LLM’s reasoning for labeling, we visualize sentiment related to OPRA concepts using tag clouds surrounding the scatter plot. These tag clouds are generated by extracting word frequencies from sentiment-encoded sentences, with word size proportional to frequency and color indicating sentiment—red for Negative and green for Positive. Highlighted words correspond to those in the sentences selected from Figure 2 (b), supporting concept-specific reasoning.
V-B Data Management View
The Data Management view, shown in Figure 2 (b), supports usability and reduces human workload. It consists of a filter (b-1) and a data table (b-2). The filter adjusts the $CoC$ range, rendering out-of-range points semi-transparent in the scatter plot (a) and hiding them from the table. The table displays sentences with $CoC$, human-coded ground truth, and LLM assessments. For a selected sentence, the clues and step-by-step reasoning for labeling are interactively linked to the view in (c). The filter identifies a subset using $CoC$, and the following operations are executed in views (a), (b), and (c) on this subset.
Before analysis, the system simplifies the dataset by removing redundant or uninformative sentences, identified via cosine distance over sentence embeddings. This reduction lowers re-evaluation costs and manual review effort. Uninformative content, such as emojis or nonsensical text (e.g., yh, hello, aaaaa, qwer), obscures OPRA concepts or adds little semantic value. The data table then supports prompt adjustment with ground-truth labels, enhancing usability by facilitating effortless problem identification and data management.
V-C LLM Reasoning & Editing View
The LLM Reasoning & Editing view in Figure 2 (c) provides an attention visualization to examine LLM’s reasoning for labeling and to support prompt modification.
Figure 5: Attention visualization in LLM Reasoning & Editing view. (a) shows the LLM input prompts divided into $S_{0\textrm{--}8}$ according to Sentence ID. (b) shows $S_{9\textrm{--}11}$, the sentences generated by LLM, and attention visualization. (b-1) is an attention visualization to identify sentences influenced by sentences generated from LLM’s assessment.
Figure 6: LLM’s reasoning for labeling on Amazon product review data. (a) shows the low $CoC$s for Satisfaction in the scatter plot. (b) presents public opinion data where LLM and experts label differently. (c) reveals that LLM has a flawed reasoning for labeling. (d) is a prompt where the user revises the clues and displays the assessment results, which show that LLM’s reasoning for labeling has been revised.
Figure 5 details the LLM Reasoning & Editing view. In (a), $S_{0\textrm{--}8}$ represent the LLM input prompts. Within this prompt structure, $S_{0\textrm{--}3}$ serve as task instructions for OPRA, while $S_{4\textrm{--}7}$ constitute an FSE ($k=1$), as described in Section IV-A. $S_{4}$ is an example input sentence, and $S_{5}$ provides the corresponding clues extracted from $S_{4}$. $S_{6}$ shows the reasoning process based on $S_{4}$ & $S_{5}$, and $S_{7}$ provides the final assessment decision. $S_{8}$ is a selected sentence from the data table in Figure 2 (b-2), which serves as the target input for the assessment. Based on this prompt, the LLM infers the label of $S_{8}$. In Figure 5 (b), the sentences $S_{9\textrm{--}11}$ are generated by the LLM according to the established reasoning pattern. $S_{9}$ identifies clues in $S_{8}$, and $S_{10}$ provides reasoning based on the clues identified in $S_{9}$. $S_{11}$ presents the final label.
Figure 5 (b-1) visualizes inter-sentence attention (ISA) [50], highlighting how the LLM utilizes preceding context. The complexity of the transformer architecture makes it unclear which elements have the greatest influence on generation. $\text{ISA}(S_{a},S_{b})$ represents the aggregated token-level attention from $S_{a}$ to $S_{b}$, and the color on each sentence dot indicates incoming attention, demonstrating how $S_{9\text{--}11}$ were influenced. In (b), $S_{10}$ is primarily influenced by $S_{0}$, which provides task instructions (e.g., “This is a satisfaction assessment tool for evaluating reviews.”). The next most influential is $S_{9}$, which is the clue sentence derived from $S_{8}$. These influences suggest that $S_{10}$ was generated appropriately.
Figure 2 (c-1) shows a pop-up window triggered by hovering over a dot in the attention visualization. It presents the generated sentence, the focused sentence, and the attention score. This view supports analysis of sentence-level influences on LLM outputs, providing insight into its reasoning process. (c) enables prompt editing based on this analysis. (c-2) includes buttons for opening the prompt editor and rerunning the LLM, with results updated in the data table in Figure 2 (b-2).
VI Use Scenarios
This section presents two real-world use scenarios demonstrating OPRA-Vis’s labeling and reasoning capabilities using Amazon [51] and Google Local [52] reviews.
Figure 7: Prompt editor. Text highlighted in orange shows the prompt before and after editing. The prompt in the blue edit box is freely editable.
VI-A Linguistic Negation Correction
This use scenario involves 198 Amazon electronics reviews [51]. The OPRA concept labels as the ground-truth of these reviews are labeled by two PR researchers. To address LLM misclassifications, our system enables analyst-driven decision-making through prompt editing. This includes certainty-based filtering, sentiment-aware tag clouds, and attention visualization. Figure 6 provides an example of the relabeling process.
Low $CoC$ often arises from linguistic phenomena such as negation, introducing ambiguity in LLM’s assessments [53]. Figure 6 (a) illustrates this issue, where OPRA-Vis indicates low $CoC$s for Satisfaction. Using the lasso selection tool, we cluster the dense blue points. The Data Management view in (b) reveals that 41 sentences fall within the $\sim0$.3 CoCCoC range, and some points are mislabeled. (b) shows a mislabeled example, “The keyboard is easy to clean, […] the clear underneath”. While the expert labels this data point as True for Satisfaction, as indicated by “Label (Expert) = True”, the LLM labels it as False. This mismatch likely happens due to LLM’s limited interpretability of natural language.
Figure 8: Correcting a sentiment-biased mislabel. (a) The LLM incorrectly labels the OPRA concept Trust as True by focusing on positive-sentiment clues, such as “fixed the problem”. (b) After PR experts rewrite few-shot examples (FSEs) in the Prompt Editor to connect Trust with cues of transparency and fairness, the LLM’s reasoning and assessment align with the expert label of False.
In Figure 6 (c), LLM misclassified the sentence “water doesn’t bother it too much”, which conveys a True meaning, as False. This is shown in CLUES (Sentence ID==9), where LLM focuses on the phrase “doesn’t bother it too much”, highlighting its difficulty in handling negation. Consequently, LLM inferred the SATISFACTION label (Sentence ID==11) as False, based on REASONING (Sentence ID==10). (c-1) visualizes that REASONING (Sentence ID==10) was most influenced by CLUES (Sentence ID==9), shown as an orange dot. For clarity, we excluded Sentence ID==0 and Sentence ID==10, both represented as red dots with the highest attention scores, from this inter-sentence attention analysis. Sentence ID==0 corresponds to Instructions and appears in nearly all generations, making it unsuitable for analyzing sentence-specific influence (see Sections V-C). Sentence ID==10 was also excluded as it is self-referential in the attention map. These findings guided prompt editing, as shown in Figure 7.
We revise the prompt as shown in Figure 7, where yellow highlights indicate the edited text. Specifically, we change “water doesn’t bother it too much” from False to True. The phrase “easy to clean” includes the word “easy”, which suggests a True connotation and appears blue in the tag cloud for “Satisfaction=True” (Figure 6 (a-1)). Therefore, this clue is retained. We also add “works great for games and office typing” to the True category in CLUES, as it supports Satisfaction. Although only the blue-highlighted text is revised in this case, any part of the prompt remains editable.
Figure 6 (d) shows the updated model assessment following the edited prompt. The CLUES at Sentence ID=9 reflects the revised input, while the REASONING at Sentence ID=10 is newly generated based on these modified clues. Influenced by this reasoning, LLM corrects its assessment of Satisfaction to True. OPRA-Vis propagates the edit by updating the shared prompt template and the exemplar pool. Then it reassesses the entire dataset, updating labels, rationales, and $CoC$ values. This process clarifies how the prompt manages phenomena like negation, updates similar cases, increases $CoC$ near relevant concept edges, and reduces inconsistencies between the LLM and expert labels.
VI-B Correcting Mislabeling Caused by Sentiment Bias
This scenario, illustrated in Figure 8, highlights a failure in concept labeling due to sentiment-bias reasoning and its correction through few-shot learning. The LLM confuses overall positivity with OPRA-informed clues, leading to the mislabeling of the sentence. We demonstrate this scenario using Google Local reviews [52], which contains 666 million reviews of U.S. businesses as of September 2021. The dataset comprises ratings, textual content, images, and metadata, including location, category, and pricing information.
Figure 8 (a) displays the initial mislabeling. In a sentence from public opinion data indicating that the issue was resolved, the LLM focuses on positive-sentiment phrases, such as “fixed the problem”. It interprets these as evidence for the OPRA concept Trust and outputs True, while the expert label is False. This error reveals a reasoning shortcut: the LLM equates problem resolution or satisfaction with trustworthiness, rather than evaluating OPRA-informed clues such as transparency, fairness, and ethical conduct. The table emphasizes the discrepancy between “Label (LLM) = True” and “Label (Expert) = False”, along with low $CoC$, which aligns with sentiment-driven reasoning, as indicated by the orange-highlighted positive sentiment in the initial LLM reasoning. Figure 8 (b) shows the correction made after PR experts rewrite and expand the FSEs in the Prompt Editor to anchor better Trust to the OPRA-informed clues, as described in Section III-A. In the corrected reasoning, the LLM explicitly dismisses positive tone as sufficient evidence and instead grounds its assessment on trustworthiness in the REASONING text. This assessment aligns with the expert label of False, demonstrating that editing FSEs to align with concepts effectively mitigates sentiment bias and restores construct-centered labeling.
VII Quantitative Evaluation
This section presents quantitative evaluations of OPRA-Vis, focusing on labeling accuracy and the effectiveness of prompts. We utilize expert-labeled datasets that include 198 Amazon product reviews (Amazon), 100 Google Local reviews (Google), and 100 Internet Movie Database reviews (IMDB[54]. Each dataset is annotated with OPRA concept labels, creating a controlled setting to compare performance across various PR domains.
VII-A Model selection via assessment accuracy
We evaluate the core model integrated into OPRA-Vis by comparing three well-established PLMs: Llama 2 (7B) [55], GPT-3 [18], and Gemma (7B) [17]. The goal of this comparison is not to establish SOTA performance, but to identify a suitable base model for OPRA-Vis within our development constraints. This ensures that the evaluation is procedural and reproducible rather than tied to the specific vintage of any given model. To ensure consistency across models, we applied CoT prompting with embedded clues and reasoning to incorporate PR theory into the evaluation process. Table I reports the labeling accuracy of concepts for each evaluated model.
TABLE I: Comparisons of OPRA concept labeling accuracy. We evaluate Llama 2 (7B), GPT-3, and Gemma (7B) on three datasets: Amazon product reviews (Amazon), Google Local reviews (Google), and Internet Movie Database reviews (IMDB). Abbreviations: Satis. = Satisfaction, Commit. = Commitment, and C.M. = Control Mutuality.
Dataset Model OPRA Labeling Accuracy Average
Trust Satis. Commit. C.M. Accuracy
One-shot settings
Amazon Llama 2 70% 79% 49% 44% 60.50%
GPT-3 79% 90% 55% 54% 69.50%
Gemma 77% 88% 63% 72% 75%
Google Llama 2 66% 70% 49% 53% 59.50%
GPT-3 76% 84% 68% 49% 69.25%
Gemma 76% 83% 72% 55% 71.50%
IMDB Llama 2 59% 70% 58% 49% 59%
GPT-3 73% 92% 72% 46% 70.75%
Gemma 75% 92% 81% 64% 78%
Few-shot ($k=8$) settings
Amazon Llama 2 67% 84% 54% 46% 62.75%
GPT-3 78% 89% 58% 55% 70%
Gemma 74% 89% 60% 87% 77.50%
Google Llama 2 70% 84% 65% 43% 65.50%
GPT-3 76% 85% 72% 46% 69.75%
Gemma 71% 84% 70% 78% 75.75%
IMDB Llama 2 64% 83% 63% 46% 64%
GPT-3 73% 94% 69% 46% 70.50%
Gemma 71% 88% 89% 97% 86.25%
The average accuracy of Gemma surpassed that of the other PLMs in both one-shot and few-shot ($k=8$) scenarios. Across all datasets, Gemma achieved the highest overall average, followed by GPT-3 and Llama 2. All models showed improved accuracy under few-shot conditions compared to one-shot, except for GPT-3, which exhibited a slight decrease in accuracy for the Google dataset. Examining individual OPRA concepts, GPT-3 achieved 1–6% higher accuracy than Gemma for Trust and Satisfaction in the one-shot Amazon dataset, as well as in few-shot evaluations across Amazon, Google, and IMDB. However, Gemma consistently outperformed GPT-3 for Commitment and Control Mutuality across most scenarios. These results indicate that OPRA-Vis effectively integrates PR theory by utilizing Gemma as its core model.
VII-B Prompting Test
[…]
Dataset Method OPRA Labeling Accuracy Average
Trust Satis. Commit. C.M. Accuracy
One-shot settings
Amazon Vanilla 71% 83% 58% 25% 59.25%
CoT 74% 86% 51% 43% 63.50%
[…]
Google Vanilla 70% 79% 66% 34% 62.25%
CoT 74% 83% 70% 39% 66.50%
[…]
IMDB Vanilla 59% 83% 74% 18% 58.50%
CoT 74% 88% 84% 32% 69.50%
[…]
Few-shot ($k=8$) settings
Amazon Vanilla 68% 83% 53% 48% 63%
CoT 70% 85% 55% 64% 68.50%
[…]
Google Vanilla 70% 82% 60% 45% 64.25%
CoT 71% 80% 66% 58% 68.75%
[…]
IMDB Vanilla 67% 58% 85% 28% 59.50%
CoT 70% 81% 84% 38% 68.25%
[…]
VIII Qualitative Evaluation
In this section, we report three qualitative evaluations: an effectiveness test, a usability test, and expert feedback. All studies were conducted under an IRB-exempt protocol approved by Sejong University, Seoul, South Korea (SUIRB-HR-E-2024-016). Informed consent was obtained from all participants prior to participation.
VIII-A Effectiveness
We recruited eight participants with backgrounds in data visualization: one undergraduate student ($PE_{0}$), one master’s student ($PE_{1}$), two doctoral students ($PE_{2,3}$), and four participants holding master’s degrees ($PE_{4\text{--}7}$). Each participant completed two tasks—manual coding and labeling using OPRA-Vis—on the same set of 20 questions (five per OPRA concept: Trust, Satisfaction, Commitment, and Control Mutuality) derived from the Amazon review dataset described in Section VI-A. For both tasks, possible responses were “True”, “False”, or “I don’t know”.
TABLE III: Comparison of OPRA labeling accuracy between human coding and our system. “I don’t know” counts appear in parentheses. Paired t-test shows $p<0.05$ (significant). P.NO. indicates participant ID.
P.NO. OPRA Labeling Accuracy Average
Trust Satis. Commit. C.M. Accuracy
Human-coding
$PE_{0}$ 100% 100% 60% (1) 100% 90% (1)
$PE_{4}$ 80% (1) 100% 40% (2) 100% 80% (3)
$PE_{1}$ 100% 100% 40% (2) 80% (1) 80% (3)
$PE_{6}$ 80% (1) 100% 20% (3) 100% 75% (4)
$PE_{2}$ 80% (1) 80% (1) 40% (3) 80% (1) 70% (6)
$PE_{3}$ 60% (1) 80% (1) 80% 100% 80% (2)
$PE_{5}$ 100% 100% 60% (1) 100% 90% (1)
$PE_{7}$ 80% (1) 80% 60% (2) 100% 80% (3)
Average 85% 92.5% 50% 95% 80.63%
Using our system
$PE_{0}$ 100% 100% 100% 80% 95%
$PE_{1}$ 100% 100% 80% 80% 90%
$PE_{4}$ 100% 100% 100% 100% 100%
$PE_{6}$ 80% 100% 100% 80% 90%
$PE_{2}$ 80% 100% 80% 80% 85%
$PE_{3}$ 100% 100% 80% (1) 100% 95% (1)
$PE_{5}$ 100% 100% 100% 80% 95%
$PE_{7}$ 100% 80% 80% 100% 90%
Average 95% 97.5% 90% 87.5% 92.5
Table III summarizes labeling accuracy. OPRA-Vis outperformed human coding for Trust, Satisfaction, and Commitment, with a 40% gain in Commitment. It also reduced “I don’t know” responses by 82.61%. A paired t-test confirmed significance ($p=0.0002<0.05$). However, accuracy for Control Mutuality decreased by 7.5% compared to human coding.
The accuracy drop was due to Control Mutuality being assessed in a review (“I honestly haven’t sought out much information about the governance of sprint. I’m sure they are willing to be transparent, I’m just not all that interested”). Participants initially labeled it False but changed to True after viewing OPRA-Vis’s reasoning. PR experts clarified that interpretation depends on whether emphasis is placed on the first or second sentence in the reviews. The experts prioritized the first sentence, reasoning that insufficient customer knowledge of company efforts undermines any basis for assessing Control Mutuality. They noted, “Uncertainty about information alone does not confirm adequate organizational effort to maintain control mutuality”. However, they acknowledged that emphasizing the second sentence could make OPRA-Vis’s decision appear reasonable. Ultimately, given linguistic nuances in PR contexts, they classified Control Mutuality as False. This discrepancy highlights a limitation of OPRA-Vis’s in-context learning, discussed in Section IX.
VIII-B Usability
We recruited 12 participants with diverse domain knowledge to evaluate OPRA-Vis usability: Mathematics, computer science, and PR ($PU_{0}$); computer science and PR ($PU_{1}$); computer science and NLP ($PU_{2,3}$); game development ($PU_{4}$); traffic and deep learning ($PU_{5}$); and computer science ($PU_{6\textrm{--}11}$). All were familiar with online reviews and experienced with data analysis tools (e.g., Excel, Tableau, IBM SPSS) and Python tool development. $PU_{0}$ and $PU_{1}$ had PR-related development experience; $PU_{0}$, $PU_{1}$, $PU_{5}$, and $PU_{6}$ had data visualization knowledge. None participated in developing OPRA-Vis. The usability test used the same dataset as in Section VII-A.
The usability test had three stages: description, utilization, and feedback. Participants first learned about PR theory, CoT prompting, and public opinion data. Then, they used OPRA-Vis for 45 minutes to explore scenarios from Section VI, adjusting labels via prompt editing. Finally, they completed a post-test system usability scale (SUS) questionnaire [56] and provided feedback. In the usability test, OPRA-Vis received an average SUS score of 84.75, with a median of 90, a minimum of 67.5, and a maximum of 95. Given that scores above 70 are generally considered acceptable [57], these results indicate that OPRA-Vis is perceived as highly usable. For each item of the SUS questionnaire ($I_{1\textrm{--}10}$), participants gave 90 ($I_{1}$), 90 ($I_{2}$), 90 ($I_{3}$), 67.5 ($I_{4}$), 90 ($I_{5}$), 77.5 ($I_{6}$), 90 ($I_{7}$), 85 ($I_{8}$), 95 ($I_{9}$), and 72.5 ($I_{10}$). Among them, $I_{4}$ (“I think that I would need the support of a technical person to be able to use this system”), $I_{6}$ (“I thought there was too much inconsistency in this system”), and $I_{10}$ (“I needed to learn a lot of things before I could get going with this system”) were evaluated as having low contribution compared to other items.
We identified reasons for low contributions on three items based on participant feedback. For item $I_{4}$, participants ($PU_{2,4,7}$) struggled due to limited understanding of data visualization. For $I_{6}$, participants ($PU_{4,7,11}$) noted inconsistencies in LLM decisions versus ground-truth, attributed to low comprehension of PR theory. For $I_{10}$, participants ($PU_{0,2,8}$) raised usability concerns: $PU_{0}$ highlighted challenges for non-PR users, $PU_{2}$ noted difficulties in writing prompts for NLP novices, and $PU_{8}$ cited limited knowledge in various areas. They suggested that a simplified OPRA-Vis could improve usability for non-experts. Positive feedback included comments such as, “It is a tool […] to easily learn and utilize PR knowledge”, and improvement suggestions included, “If there were a system that showed the difference between pre-feedback and post-feedback, […] feedback would be clearer”.
VIII-C Expert Feedback
Seven experts evaluated OPRA-Vis’s effectiveness and suggested improvements. Two specialize in PR theory ($E_{0}$, $E_{1}$), one in PR-based deep learning model development ($E_{2}$), one in PR-based visual analysis ($E_{3}$), two in NLP ($E_{4}$, $E_{5}$), and one in data visualization ($E_{6}$). $E_{1}$ and $E_{2}$ contributed to OPRA-Vis’s design and ground truth generation; others encountered OPRA-Vis for the first time after development.
Effectiveness: Experts praised the intuitive visualizations. In tag clouds, common words or company names such as “type, five, trip, Amazon, Apple” are typically excluded from sentiment analysis because they lack inherent semantic meaning. However, in public opinion data, these words can carry significance. For example, in the tag cloud for “Trust=True” shown in Figure 6 (a), “Amazon” exhibits positive sentiment, as it frequently appears in contexts like “I order a lot of stuff from Amazon. They always deliver on time”, which provides True clues for Trust. $E_{0}$ noted, “The bipolar illustration of key relational dimensions like commitment and trust allows PR managers to perceive and articulate the strength and tendencies in data distributions of relationships”. $E_{6}$ remarked, “This view aids in understanding overall opinion distributions and common keywords. It serves as a starting point for selecting relevant comments to verify if the LLM operated as intended”. Both $E_{3}$ and $E_{6}$ highlighted the system’s effectiveness: “This view promptly helps understand how LLM evaluates each comment. The displayed texts of clues and reasoning are useful for understanding how LLM labeled the comment. […]; In particular […], attention visualization greatly assists reasoning”. Moreover, $E_{1}$ emphasized, “[…] it is important to note that human coding can often be inconsistent and not always accurate”, and concluded that OPRA-Vis successfully addresses the limitations of existing human coding.
Improvements: Experts suggested extending OPRA-Vis to include time-series analysis of public opinion, beyond its current focus on reasoning and correction. $E_{0}$ noted, “Time-series data is also critical. […] Real-time tracking enables managers to observe historical changes in relationships. […] can inform PR managers about events triggering changes (True or False) and allow for the preparation of appropriate managerial and communication responses”. $E_{5}$ recommended adding automation and visual reliability cues, stating, “Wouldn’t it be possible to approach things in a more user-friendly way by leaving the variables untouchable and, if there is nothing to use, pressing None, as LLM solves the problem? […] I think that if the reliability of the LLM is drawn in a score or graph, we can rely on the LLM and use it”. Future work will focus on generating labeled data from OPRA-Vis to enable fine-tuning and exploring visual analysis techniques for tracking opinion trends over time.
IX Conclusion and Discussion
This paper introduced OPRA-Vis, a visual analytics system supporting organization-public relationship assessment (OPRA) by leveraging LLMs and CoT prompting for expert-driven labeling. Integrating PR’s OPRA concepts, it semi-automates public opinion labeling with CoT prompting enriched by expert-informed clues and reasoning. OPRA-Vis addresses the linguistic challenges and limitations of CoT prompting by visualizing LLM’s reasoning for labeling using the certainty of concepts ($CoC$), sentiment, and attention scores. The system supports interactive analysis using the OPRA Concept Space view, Data Management view, and LLM Reasoning & Editing view. Evaluations, which include use scenarios, quantitative tests, usability studies, and expert feedback, demonstrate an effective integration of expert knowledge with LLM reasoning without the need for fine-tuning, although some limitations remain.
In-context learning: In-context learning through CoT prompting with few-shot examples (FSEs) is widely used and can be useful for reasoning tasks [18], but often underperforms fine-tuned models in text assessment due to language complexity [15]. Its effectiveness heavily depends on the structure and content of prompts, requiring domain expertise [14]. However, domain experts often need assistance with LLM-specific techniques. To address linguistic challenges and improve reliability, we integrated visual analysis to clarify LLM reasoning. We incorporated expert knowledge by defining OPRA concepts, including ten definitions of top-level and sub-dimensions [13]. To ensure clarity and relevance, we adopted few-shot CoT prompting with clues and reasons [15]. Still, further work is needed to improve reasoning in in-context learning, and to address the constraints of current fine-tuning techniques, e.g., parameter-efficient tuning [58].
Figure 9: Application of OPRA-Vis to Jigsaw toxic comments dataset.
Domain applicability: While OPRA-Vis is designed to address challenges in the PR domain, its framework can be adapted to other domains. As shown in Figure 9, we applied OPRA-Vis to the Jigsaw toxic comments dataset [59], which has five dimensions. Datasets with more than five dimensions challenge the scatter plot with tag clouds, as True and False labels are arranged symmetrically ($5\times 2$ sides). With more sides, the layout becomes circular, making it harder to interpret relationships. Future work will develop a visualization to handle multiple dimensions effectively.
Acknowledgment
This should be a simple paragraph before the bibliography to thank those individuals and institutions who have supported your work on this article.

### v2

@@
 These metrics are typically obtained via surveys, supporting relationship improvement strategies across sectors such as business and government, and helping organizations strengthen public engagement over both short- and long-term horizons.
-PR analysis involves two distinct classification tasks: the binary assessment of OPRA concepts (labeling opinions as True or False for each concept) and sentiment analysis, which classifies the emotional tone of public opinion as Positive or Negative.
+In PR analysis, OPRA concepts are evaluated at the sentence level.
+A sentence is labeled True for a concept if it contains identifiable clues supporting that concept; otherwise, it is labeled False.
+These binary labels denote the presence or absence of a concept, rather than the overall quality of the relationship.
+PR experts rely on cited clues and the distribution of True/False sentences to inform policy recommendations for strengthening organization–public relationships.
+Sentiment analysis complements OPRA labeling by capturing the emotional tone of a sentence (e.g., Positive or Negative), which is not conveyed by concept labels.
 III-B Expert Requirements
@@
     The system enhances usability through task-relevant filtering informed by OPRA labeling (C2, C3, R6).
-Integrating PR theory with LLMs for labeling public opinion data faces three main challenges (C1–C3) as introduced in Section I.
-To address the challenge of insufficient training data for fine-tuning LLMs (G1), we implement the Gemma [17] model with CoT prompting, based on in-context learning.
-This method leverages textitFSEs provided by PR experts to label the data.
+Integrating PR theory with LLMs for labeling public opinion data faces main challenges (C1–C3) as introduced in Section I.
+To address the challenge of insufficient training data for fine-tuning LLMs (G1), we implement the Gemma [17] with CoT prompting, based on in-context learning.
+This method leverages FSEs provided by PR experts to label the data.
 Moreover, we have developed an LLM pipeline that represents OPRA-related clues in public opinion data as a continuous encoding (G2).
-Adopting a BERT [19] model reduces the dependence on human effort to classify sentiment by analyzing public opinion sentences.
+Adopting a BERT [19] reduces the dependence on human effort to classify sentiment by analyzing public opinion sentences.
 Sentiment classification enriches labeled data and broadens the analyses supported by our system (G2).
-We combine visual analytics with CoT prompting that incorporates clues and reasons to audit and revise the model’s reasoning for labeling, supported by coordinated views and prompt editing (G3, G4, G5).
+We combine VA with CoT prompting that incorporates clues and reasons to audit and revise the model’s reasoning for labeling, supported by coordinated views and prompt editing (G3, G4, G5).
 CoT prompting elicits “think-aloud” rationales, making the LLM’s reasoning for labeling explicit and auditable (G3).
@@
 The Data Management view features a slider for data filtering, as presented in (b-1).
-The filter uses the CoCCoC range to update point visibility in the scatter plot in (a).
+The filter uses the $CoC$ range to update point visibility in the scatter plot in (a).
 The data table includes certainty, concept labels from LLM and experts, and data exclusion, as shown in (b-2).
@@
 Here, the max operation selects the highest attention score within each expert instruction, while the mean operation averages these maximum scores across multiple expert instructions for the same concept-label pair.
-The computed $CoC$ values are standardized and rescaled to $[0,1]$ for consistent comparison and intuitive interpretation: values near $1$ indicate high certainty for True OPRA concepts, values near 0 indicate high certainty for False OPRA concepts, and $0.5$ denotes neutral certainty.
+The computed $CoC$ values are standardized and rescaled to $[0,1]$ for consistent comparison and intuitive interpretation.
 These $CoC$ values serve as quantitative indicators of concept certainty, supporting analytical evaluation.
@@
 (a) Before: Initial t-SNE layout of sentence embeddings forms dense clusters without clear concept alignment.
-(b) After: A gravity model uses True and False concept vertices ($\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ in Equation 2) as gravitational attractors, where each sentence is pulled toward its corresponding concept edge with force proportional to its certainty (CoCCoC).
-High-certainty sentences experience a stronger gravitational pull and move closer to concept edges.
-In contrast, low-certainty sentences experience weaker forces and remain near the center, resulting in eight radial structures separating concepts and certainty levels.
+(b) After: A gravity model uses True and False label vertices ($\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ in Equation 2) as gravitational attractors, where each sentence is pulled toward its corresponding concept edge with force proportional to its CoCCoC deviation from 0.5.
+Sentences with CoCCoC values distant from 0.5 are pulled more strongly toward concept edges, while those near 0.5 experience weaker forces and remain closer to the center.
+This produces eight radial structures that separate concepts by certainty levels.
 The OPRA Concept Space view, shown in Figure 2 (a), visualizes public opinion texts according to the four OPRA concepts—Trust, Satisfaction, Commitment, and Control Mutuality—using a scatter plot and tag clouds.
@@
 To better reflect OPRA-specific relationships, we further adjust the positions using Equation 2, which incorporates the $CoC$.
-The gravitational adjustment ensures that as $CoC$ increases, points move toward the corresponding concept edge.
-To achieve this adjustment, we employ a physics-inspired gravity model [49], where the vertices representing OPRA concepts act as gravitational attractors.
+We employ a physics-inspired gravity model [49], where the vertices representing OPRA concepts act as gravitational attractors.
 Sentence positions are iteratively updated based on their $CoC$ values through a constrained simulation.
@@
 $\mathbf{u}_{i}^{(t+1)}$ $=\gamma\cdot\mathbf{u}_{i}^{(t)}+\delta\cdot\sum_{j}\mathbf{F}_{ij};\quad\mathbf{p}_{i}^{(t+1)}=\mathbf{p}_{i}^{(t)}+\mathbf{u}_{i}^{(t+1)}$ (2)
-Here, $\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ represent True and False concept vertices for concept $j$, $\mathbf{1}_{\{\cdot\}}$ is the indicator function, $\alpha_{\text{base}}$ is the base gravitational strength, $G$ is the gravitational constant, $r_{ij}=\|\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}\|_{2}$ is the Euclidean distance, $\mathbf{u}_{i}^{(t)}$ is the velocity vector, $\gamma$ is the damping factor, $\delta$ is the step size, $\epsilon_{1}$ prevents force divergence when $r_{ij}=0$, and $\epsilon_{2}$ prevents division by zero when computing directions.
+Here, $\mathbf{v}_{j}^{+}$ and $\mathbf{v}_{j}^{-}$ represent True and False label vertices for concept $j$, $\mathbf{1}_{\{\cdot\}}$ is the indicator function, $\alpha_{\text{base}}$ is the base gravitational strength, $G$ is the gravitational constant, $r_{ij}=\|\mathbf{v}_{\text{target}}^{(i,j)}-\mathbf{p}_{i}^{(t)}\|_{2}$ is the Euclidean distance, $\mathbf{u}_{i}^{(t)}$ is the velocity vector, $\gamma$ is the damping factor, $\delta$ is the step size, $\epsilon_{1}$ prevents force divergence when $r_{ij}=0$, and $\epsilon_{2}$ prevents division by zero when computing directions.
 In our implementation, we use $\alpha_{\text{base}}=2.0$, $G=1.0$, $\gamma=0.8$, $\delta=0.1$, $\epsilon_{1}=0.01$, and $\epsilon_{2}=10^{-10}$.
@@
 We will explore systematic optimization methods in future work.
-The gravitational force follows an inverse square law, where $\alpha_{ij}$ acts as the gravitational mass proportional to concept certainty.
-When $CoC_{ij}=0.5$, no gravitational force is applied as the sentence shows neutral certainty.
+The gravitational force follows an inverse square law, where $\alpha_{ij}$ acts as the gravitational mass proportional to the deviation of $CoC_{ij}$ from 0.5.
+When $CoC_{ij}~=~0.5$, no gravitational force is applied as the sentence shows intermediate $CoC$ values.
 Positions are constrained within a unit circle through normalization to maintain visual coherence.
-This iterative gravitational simulation continues for up to 200 iterations or until position changes fall below $10^{-4}$, creating spatial clustering where sentences are gravitationally attracted toward vertices corresponding to their respective concepts.
+This iterative gravitational simulation continues for up to 200 iterations or until position changes fall below $10^{-4}$, creating spatial clustering where sentences are gravitationally attracted toward vertices corresponding to their respective concept labels.
 Figure 3 illustrates this transformation: (a) shows the initial t-SNE layout, and (b) shows the adjusted plot reflecting $CoC$ values with gravitational attractors.
@@
 (a) to (d) show the views with Trust, Satisfaction, Commitment, and Control Mutuality selected, respectively.
-The scatter plot enables interactive concept selection: clicking on an OPRA concept label on the concept edge updates the color encoding based on the $CoC$ values for that concept.
-We use a blue-yellow-red color scheme, where blue indicates low certainty (near 0), yellow represents neutral certainty (around 0.5), and red corresponds to high certainty (near 1).
-This three-color gradient aligns naturally with the $CoC$ scale, which is centered around the neutral value of 0.5.
+The scatter plot enables interactive concept selection.
+Clicking on an OPRA concept label on the concept edge updates the color encoding based on the CoC values for that concept using a blue-yellow-red color scheme.
 Figures 4 (a)–(d) illustrate scatter plots for the cases when the user selects Trust, Satisfaction, Commitment, and Control Mutuality, respectively.
-To highlight regions of high and low certainty, we provide a distribution chart for the user-selected OPRA concept label from the scatter plot, as illustrated in Figure 2 (a-1).
-The data are aligned along an axis whose endpoints correspond to the concept’s True and False values.
-Each point is orthogonally projected onto this axis, producing a position in $[0,1]$ based on its $CoC$.
-The distribution is shown as a histogram with a $\log_{10}$-scaled y-axis to emphasize patterns across frequency ranges.
-Users can switch between linear, $\ln$, $\log_{2}$, and $\log_{10}$ scales by clicking the chart title, allowing flexible analysis of both sparse and dense regions.
+To illustrate how $CoC$ values translate into positions along the concept axis, we provide a distribution chart for the user-selected OPRA concept label from the scatter plot, as illustrated in Figure 2 (a-1).
+The axis spans from False to True, and each data point is orthogonally projected onto the axis, yielding a position within $[0,1]$.
+The distribution is shown as a histogram with bars stacked by $CoC$ values and a $\log_{10}$-scaled y-axis to emphasize patterns across frequency ranges.
+Users can toggle between linear, $\ln$, $\log_{2}$, and $\log_{10}$ scales by clicking on the chart title, enabling flexible analysis of both sparse and dense regions of the position distribution.
 To facilitate analysis of LLM’s reasoning for labeling, we visualize sentiment related to OPRA concepts using tag clouds surrounding the scatter plot.
@@
 Figure 6: LLM’s reasoning for labeling on Amazon product review data.
-(a) shows the low $CoC$s for Satisfaction in the scatter plot.
-(b) presents public opinion data where LLM and experts label differently.
+(a) shows the OPRA Scpace view with Satisfaction selected by the user.
+b) presents public opinion data where LLM and experts label differently.
 (c) reveals that LLM has a flawed reasoning for labeling.
@@
 VI-A Linguistic Negation Correction
-This use scenario involves 198 Amazon electronics reviews [51].
+This use scenario involves 198 Amazon product reviews [51].
 The OPRA concept labels as the ground-truth of these reviews are labeled by two PR researchers.
@@
 Figure 6 provides an example of the relabeling process.
-Low $CoC$ often arises from linguistic phenomena such as negation, introducing ambiguity in LLM’s assessments [53].
-Figure 6 (a) illustrates this issue, where OPRA-Vis indicates low $CoC$s for Satisfaction.
-Using the lasso selection tool, we cluster the dense blue points.
-The Data Management view in (b) reveals that 41 sentences fall within the $\sim0$.3 CoCCoC range, and some points are mislabeled.
-(b) shows a mislabeled example, “The keyboard is easy to clean, […] the clear underneath”.
-While the expert labels this data point as True for Satisfaction, as indicated by “Label (Expert) = True”, the LLM labels it as False.
-This mismatch likely happens due to LLM’s limited interpretability of natural language.
+Guided by the distribution chart in Figure 6 (a), we partition the data into three $CoC$ groups, as shown within the three gray dotted frames.
+While exploring the group (a-2) with $CoC$ in the range of 0 to 0.3 using the filtering on the Data Management view in (b), we capture that 41 sentences surface several mismatches between expert and LLM labels.
+(b) shows a representative case, “The keyboard is easy to clean, […] the clear underneath”.
+The expert labels this sentence True for Satisfaction, as indicated by “Label (Expert) = True”, whereas the LLM labels it False.
+As we examine the LLM’s reasoning for labeling in (c), we find that linguistic negation underlies this error.
+Negation is a well-known source of ambiguity in LLM’s labeling tasks [53].
 Figure 8: Correcting a sentiment-biased mislabel.
@@
 For clarity, we excluded Sentence ID==0 and Sentence ID==10, both represented as red dots with the highest attention scores, from this inter-sentence attention analysis.
-Sentence ID==0 corresponds to Instructions and appears in nearly all generations, making it unsuitable for analyzing sentence-specific influence (see Sections V-C).
+Sentence ID==0 corresponds to Instructions and appears in nearly all generations, making it unsuitable for analyzing sentence-specific influence (see Section V-C).
 Sentence ID==10 was also excluded as it is self-referential in the attention map.
@@
 OPRA-Vis propagates the edit by updating the shared prompt template and the exemplar pool.
-Then it reassesses the entire dataset, updating labels, rationales, and $CoC$ values.
-This process clarifies how the prompt manages phenomena like negation, updates similar cases, increases $CoC$ near relevant concept edges, and reduces inconsistencies between the LLM and expert labels.
+Then it reassesses the entire dataset, updating labels and rationales.
+This process clarifies how the prompt manages phenomena like negation, updates similar cases, and reduces inconsistencies between the LLM and expert labels.
 VI-B Correcting Mislabeling Caused by Sentiment Bias
@@
 This error reveals a reasoning shortcut: the LLM equates problem resolution or satisfaction with trustworthiness, rather than evaluating OPRA-informed clues such as transparency, fairness, and ethical conduct.
-The table emphasizes the discrepancy between “Label (LLM) = True” and “Label (Expert) = False”, along with low $CoC$, which aligns with sentiment-driven reasoning, as indicated by the orange-highlighted positive sentiment in the initial LLM reasoning.
+The table emphasizes the discrepancy between “Label (LLM) = True” and “Label (Expert) = False”, which aligns with sentiment-driven reasoning, as indicated by the orange-highlighted positive sentiment in the initial LLM reasoning.
 Figure 8 (b) shows the correction made after PR experts rewrite and expand the FSEs in the Prompt Editor to anchor better Trust to the OPRA-informed clues, as described in Section III-A.